# Animate Your Thoughts: Decoupled Reconstruction of Dynamic Natural Vision from Slow Brain Activity

### v1

@@ C.4 Algonauts2021
This dataset is publicly released for the 2021 Algonauts Challenge[49]. During data acquisition, 10 participants passively view 1100 silent videos of everyday events, each approximately 3 seconds in duration, presented three times. The participants’ fMRI is recorded using a 3T Trio Siemens scanner with a spatial resolution of 2.5 millimeters and a repetition time (TR) of 1.75 seconds. The fMRI preprocessing involves steps such as slice-timing correction, realignment, coregistration, and normalization to the MNI space. Additionally, the fMRI data are interpolated to a TR of 2 seconds.
@@ C.5 Data Acquisition
Appendix D Implementation Details
D.1 Hyperparameter Settings
For all three datasets employed in the experiments, during the training of the Semantic Decoder, we set $\alpha$ to 0.5, $\lambda_{1}$ to 0.01, and $\lambda_{2}$ to 0.5. The batch size is set to 64, and the learning rate is set to 2e-4, with training conducted 100 epochs. Given the critical role of data volume and augmentation methods in the training of contrastive learning models, and the scarcity of video-fMRI paired data, we implement specific data augmentation techniques to prevent overfitting. For fMRI data, we randomly select 20% of the voxels during each iteration, zero out 50% of their values. For image data, we randomly crop one frame from eight video frames to a size of 400x400 pixels, and then resize it to 224x224. For text data, due to the presence of similar video clips in the training set (derived from a complete video segment), BLIP2 often provide identical textual descriptions, which leads to overfitting. To mitigate this, we apply more aggressive augmentation techniques. For an input sentence, we perform Synonym Replacement with a 50% probability and Random Insertion, Random Swap, and Random Deletion with a 20% probability each.
@@ D.1 Hyperparameter Settings
When training the Consistency Motion Generator, we set the mask ratio of the Sparse Causal mask to 0.6 during the training phase, with a batch size of 64 and a learning rate of 4e-5. Similarly, for stability, we implement a 50-step warmup followed by 300 epochs of training.
@@
According to Figure 8, in the inference phase, we utilize model parameters saved at the 30th, 75th, and 70th epochs for the Semantic Decoder, Structural Decoder, and Consistency Motion Generator, respectively. For the generative model, we employ inflated Stable Diffusion V1-5. Given that Stable Diffusion operates on Gaussian-distributed latent space inputs (i.e., $Z_{T}$) for Text-to-Image task, and the distribution of decoded structural information does not align with this, we apply 250 steps of Gaussian smoothing to it, which includes 50 steps of ddim inversion. All experiments are conducted on a single A100 80G GPU, with the training phase taking 8 hours and the inference phase taking 12 hours for each dataset.
D.2 Evaluation Metric Implementation
@@
This algorithm performs the N-trial n-way top-1 classification test. We describe our evaluation method in Algorithm 2. For the image-based scenario, we utilize a Vision Transformer (ViT)[50] pre-trained on ImageNet as the classifier. For video-based scenario, a pre-trained VideoMAE[52] is employed as the classifier.
@@
Spaciotemporal-level
Considering the input to the video reconstruction task contains substantial noise, there are instances where every pixel of each reconstructed video frame is either zero or noise, which would artificially inflate the CLIP score if used directly. Therefore, we calculate the CLIP score only when the VIFI-CLIP value exceeds 0.6; otherwise, we assign a score of 0.
@@ Appendix E Additional Experimental Results
E.1 Further Results on the CC2017 Dataset
As depicted in Table 5 and 6, we further demonstrate the experimental results on subjects 2 and 3 from the CC2017 dataset. Specifically, for the metrics 2-way-I and 2-way-V, we present the mean and standard deviation across all samples over 100 trials, for the remaining metrics, we offer the mean and standard deviation of all results within the test set to reflect the overall performance in reconstruction.
E.1.1 CC2017 sub2
As shown in Table 5, our model on subject 2 achieves SOTA performance in five out of seven metrics and secured second place in one.
Semantic-level $\uparrow$ Pixel-level $\uparrow$ ST-level $\uparrow$
2-way-I 2-way-V VIFI-score SSIM PSNR Hue-pcc CLIP-pcc
Nishimoto[28] 0.787±$\pm 0.04$ —— —— 0.112±$\pm 0.07$ 8.592±$\pm 2.11$ 0.713±$\pm 0.11$ ——
Wen[31] 0.783±$\pm 0.03$ —— —— $\pm 0.14$ 8.415±$\pm 3.46$ 0.626±$\pm 0.12$ ——
Wang[32] 0.727±$\pm 0.04$ 0.779±$\pm 0.03$ 0.596±$\pm 0.07$ 0.107±$\pm 0.06$ 10.940±$\pm 1.83$ 0.589±$\pm 0.19$ 0.404±$\pm 0.38$
Kupershmidt[34] 0.776±$\pm 0.03$ 0.766±$\pm 0.03$ 0.591±$\pm 0.08$ 0.157±$\pm 0.09$ 11.914±$\pm 1.89$ 0.601±$\pm 0.18$ 0.382±$\pm 0.41$
Chen[36] 0.789±$\pm 0.03$ 0.842±$\pm 0.03$ 0.595±$\pm 0.08$ 0.172±$\pm 0.09$ 8.929±$\pm 1.62$ 0.773±$\pm 1.13$ 0.409±$\pm 0.46$
Ours(sub2) 0.804±$\pm 0.03$ 0.832±$\pm 0.03$ 0.604±$\pm 0.08$ 0.287±$\pm 0.11$ 9.049±$\pm 1.45$ 0.795±$\pm 0.12$ $\pm 0.42$
Table 5: Quantitative comparison of reconstruction results on subjects 2.
E.1.2 CC2017 sub3
As shown in Table 6, our model on subject 3 achieves SOTA performance in four out of seven metrics and secured second place in two. It should be noted that the primary objective of this work is to attempt the disentanglement of semantic, structural, and motion information from fMRI data, rather than achieving SOTA performance. The superior results obtained in the aforementioned metrics can more compellingly demonstrate the effectiveness of our proposed model.
Semantic-level $\uparrow$ Pixel-level $\uparrow$ ST-level $\uparrow$
2-way-I 2-way-V VIFI-score SSIM PSNR Hue-pcc CLIP-pcc
Nishimoto[28] 0.712±$\pm 0.04$ —— —— 0.128±$\pm 0.07$ 8.546±$\pm 1.91$ 0.746±$\pm 0.12$ ——
Wen[31] —— —— —— —— —— —— ——
Wang[32] 0.722±$\pm 0.04$ 0.778±$\pm 0.03$ 0.584±$\pm 0.07$ 0.098±$\pm 0.06$ 10.758±$\pm 2.00$ 0.572±$\pm 0.49$ 0.392±$\pm 0.41$
Kupershmidt[34] 0.767±$\pm 0.03$ 0.766±$\pm 0.03$ 0.597±$\pm 0.08$ 0.128±$\pm 0.08$ 11.237±$\pm 1.86$ 0.641±$\pm 0.14$ $\pm 0.37$
Chen[36] 0.811±$\pm 0.03$ 0.848±$\pm 0.03$ 0.597±$\pm 0.08$ 0.187±$\pm 0.10$ 9.013±$\pm 1.66$ 0.771±$\pm 0.13$ 0.410±$\pm 0.46$
Ours(sub3) 0.792±$\pm 0.03$ 0.833±$\pm 0.02$ 0.600±$\pm 0.07$ 0.349±$\pm 0.11$ 9.306±$\pm 1.54$ 0.791±$\pm 0.12$ 0.415±$\pm 0.39$
Table 6: Quantitative comparison of reconstruction results on subjects 3.
E.1.3 More Reconstruction Results on Multiple Subjects
Figure 10: More reconstruction results on the CC2017 dataset.
Figure 11: Reconstruction failure cases.
E.2 Further Results on the HCP Dataset
To validate the broad effectiveness of our proposed model across different datasets, we also conduct experiments on the HCP dataset. The selection of three subjects (100610, 102816, and 104416) aligns with that of Wang et al.[32], and the parameter settings for Mind-Animator during both training and testing are identical to those used on the CC2017 dataset. We replicate Nishimoto et al.[28] ’s model on HCP, utilizing video clips from the training sets of both CC2017 and HCP as video priors. Wang only reports SSIM and PSNR metrics on the HCP dataset. Visual comparisons are presented in Figure 12, while quantitative comparisons are detailed in Table 7. The results indicate that our model still performs well across multiple subjects in the HCP dataset.
Sub ID Semantic-level $\uparrow$ Pixel-level $\uparrow$ ST-level $\uparrow$
2-way-I 2-way-V VIFI-score SSIM PSNR Hue-pcc CLIP-pcc
Sub01 Nishimoto[28] 0.658±$\pm 0.04$ —— —— 0.307±$\pm 0.14$ 11.711±$\pm 2.12$ 0.649±$\pm 0.12$ ——
Wang[32] —— —— —— 0.154±$\pm 0.11$ 13.200±$\pm 4.01$ —— ——
Ours 0.819±$\pm 0.03$ 0.783±$\pm 0.03$ 0.613±$\pm 0.07$ 0.325±$\pm 0.10$ 10.757±$\pm 2.40$ 0.820±$\pm 0.13$ 0.476±$\pm 0.41$
Sub02 Nishimoto[28] 0.661±$\pm 0.04$ —— —— 0.338±$\pm 0.14$ 11.249±$\pm 2.17$ 0.643±$\pm 0.12$ ——
Wang[32] —— —— —— 0.178±$\pm 0.11$ 13.700±$\pm 5.00$ —— ——
Ours 0.756±$\pm 0.03$ 0.759±$\pm 0.04$ 0.609±$\pm 0.07$ 0.371±$\pm 0.12$ 11.894±$\pm 2.66$ 0.834±$\pm 0.14$ 0.485±$\pm 0.44$
Sub03 Nishimoto[28] 0.657±$\pm 0.04$ —— —— 0.319±$\pm 0.14$ 10.988±$\pm 2.14$ 0.645±$\pm 0.13$ ——
Wang[32] —— —— —— 0.147±$\pm 0.11$ 12.200±$\pm 4.00$ —— ——
Ours 0.781±$\pm 0.03$ 0.793±$\pm 0.03$ 0.634±$\pm 0.07$ 0.336±$\pm 0.12$ 11.018±$\pm 2.62$ 0.834±$\pm 0.14$ 0.573±$\pm 0.39$
Table 7: Quantitative comparison of reconstruction results on HCP dataset.
E.3 Further Results on the Algonauts2021 Dataset
Sub ID Semantic-level $\uparrow$ Pixel-level $\uparrow$ ST-level $\uparrow$
2-way-I 2-way-V VIFI-score SSIM PSNR Hue-pcc CLIP-pcc
Sub01 Nishimoto[28] 0.688±$\pm 0.04$ —— —— 0.446±$\pm 0.08$ 9.626±$\pm 1.61$ 0.672±$\pm 0.15$ ——
Ours 0.722±$\pm 0.04$ 0.790±$\pm 0.03$ 0.599±$\pm 0.07$ 0.401±$\pm 0.10$ 10.088±$\pm 1.56$ 0.824±$\pm 0.09$ 0.439±$\pm 0.42$
Sub02 Nishimoto[28] 0.682±$\pm 0.04$ —— —— 0.443±$\pm 0.09$ 9.553±$\pm 1.53$ 0.676±$\pm 0.13$ ——
Ours 0.734±$\pm 0.04$ 0.765±$\pm 0.03$ 0.596±$\pm 0.07$ 0.465±$\pm 0.11$ 10.932±$\pm 1.68$ 0.796±$\pm 0.12$ 0.425±$\pm 0.43$
Sub03 Nishimoto[28] 0.701±$\pm 0.04$ —— —— 0.441±$\pm 0.08$ 9.576±$\pm 1.74$ 0.682±$\pm 0.16$ ——
Ours 0.679±$\pm 0.04$ 0.794±$\pm 0.03$ 0.591±$\pm 0.07$ 0.466±$\pm 0.12$ 11.089±$\pm 1.65$ 0.863±$\pm 0.11$ 0.397±$\pm 0.43$
Sub04 Nishimoto[28] 0.702±$\pm 0.04$ —— —— 0.446±$\pm 0.09$ 9.537±$\pm 1.61$ 0.665±$\pm 0.16$ ——
Ours 0.673±$\pm 0.04$ 0.810±$\pm 0.03$ 0.587±$\pm 0.07$ 0.479±$\pm 0.13$ 11.410±$\pm 1.81$ 0.848±$\pm 0.11$ 0.381±$\pm 0.42$
Sub05 Nishimoto[28] 0.676±$\pm 0.04$ —— —— 0.442±$\pm 0.08$ 9.498±$\pm 1.39$ 0.650±$\pm 0.14$ ——
Ours 0.689±$\pm 0.04$ 0.810±$\pm 0.03$ 0.592±$\pm 0.07$ 0.458±$\pm 0.12$ 10.814±$\pm 1.60$ 0.807±$\pm 0.12$ 0.406±$\pm 0.44$
Sub06 Nishimoto[28] 0.694±$\pm 0.04$ —— —— 0.444±$\pm 0.08$ 9.526±$\pm 1.57$ 0.665±$\pm 0.14$ ——
Ours 0.709±$\pm 0.04$ 0.783±$\pm 0.03$ 0.597±$\pm 0.06$ 0.489±$\pm 0.13$ 11.337±$\pm 1.70$ 0.834±$\pm 0.12$ 0.446±$\pm 0.43$
Sub07 Nishimoto[28] 0.674±$\pm 0.04$ —— —— 0.446±$\pm 0.08$ 9.630±$\pm 1.57$ 0.672±$\pm 0.14$ ——
Ours 0.681±$\pm 0.04$ 0.802±$\pm 0.03$ 0.578±$\pm 0.07$ 0.458±$\pm 0.11$ 10.889±$\pm 1.58$ 0.857±$\pm 0.11$ 0.329±$\pm 0.42$
Sub08 Nishimoto[28] 0.696±$\pm 0.04$ —— —— 0.444±$\pm 0.08$ 9.664±$\pm 1.57$ 0.662±$\pm 0.15$ ——
Ours 0.709±$\pm 0.04$ 0.802±$\pm 0.03$ 0.592±$\pm 0.07$ 0.467±$\pm 0.13$ 10.893±$\pm 1.56$ 0.820±$\pm 0.13$ 0.376±$\pm 0.43$
Sub09 Nishimoto[28] 0.673±$\pm 0.04$ —— —— 0.445±$\pm 0.08$ 9.573±$\pm 1.67$ 0.661±$\pm 0.14$ ——
Ours 0.731±$\pm 0.04$ 0.788±$\pm 0.03$ 0.594±$\pm 0.06$ 0.502±$\pm 0.12$ 11.310±$\pm 1.57$ 0.820±$\pm 0.12$ 0.400±$\pm 0.42$
Sub10 Nishimoto[28] 0.685±$\pm 0.04$ —— —— 0.441±$\pm 0.08$ 9.598±$\pm 1.79$ 0.662±$\pm 0.14$ ——
Ours 0.684±$\pm 0.04$ 0.777±$\pm 0.03$ 0.590±$\pm 0.07$ 0.465±$\pm 0.13$ 11.128±$\pm 1.91$ 0.858±$\pm 0.12$ 0.408±$\pm 0.43$
Table 8: Quantitative comparison of reconstruction results on Algonauts2021 dataset.
E.4 A Detailed Ablation Study on Various Hyperparameters
@@
0.787±$\pm 0.03$ 0.812±$\pm 0.03$ 0.584±$\pm 0.08$
Table 9: Ablation study on multimodal contrastive learning. Colors reflect statistical significance (paired t-test) compared to the Full Model. $p<0.0001$ (purple); $p<0.01$ (pink); $p<0.05$ (yellow); $p>0.05$ (green).
@@
Pixel-$\uparrow$
SSIM PSNR Hue-pcc
Ratio=0 0.297±$\pm 0.11$ 9.037±$\pm 1.42$ 0.758±$\pm 0.11$
Ratio=0.2 0.276±$\pm 0.11$ 8.847±$\pm 1.59$ 0.767±$\pm 0.12$
Ratio=0.4 0.285±$\pm 0.10$ 9.045±$\pm 1.48$ 0.768±$\pm 0.11$

### v2

@@ B.4 Algonauts2021
This dataset is publicly released for the 2021 Algonauts Challenge (Cichy et al. (2021)). During data acquisition, 10 participants passively view 1100 silent videos of everyday events, each approximately 3 seconds in duration, presented three times. The participants’ fMRI is recorded using a 3T Trio Siemens scanner with a spatial resolution of 2.5 millimeters and a repetition time (TR) of 1.75 seconds. The fMRI preprocessing involves steps such as slice-timing correction, realignment, coregistration, and normalization to the MNI space. Additionally, the fMRI data are interpolated to a TR of 2 seconds.
@@ B.5 Data Acquisition
Appendix C Implementation Details
C.1 Hyperparameter Settings
For all three datasets employed in the experiments, during the training of the Semantic Decoder, we set $\alpha$ to 0.5, $\lambda_{1}$ to 0.01, and $\lambda_{2}$ to 0.5. The batch size is set to 64, and the learning rate is set to 2e-4, with training conducted 100 epochs. Given the critical role of data volume and augmentation methods in the training of contrastive learning models, and the scarcity of video-fMRI paired data, we implement specific data augmentation techniques to prevent overfitting. For fMRI data, we randomly select 20% of the voxels during each iteration, zero out 50% of their values. For image data, we randomly crop one frame from eight video frames to a size of 400x400 pixels, and then resize it to 224x224. When extracting the CLIP image features $\mathbf{v}$ for each video, we input each frame into the CLIP visual encoder and then compute the average across all frames. For text data, due to the presence of similar video clips in the training set (derived from a complete video segment), BLIP2 often provide identical textual descriptions, which leads to overfitting. To mitigate this, we apply more aggressive augmentation techniques. For an input sentence, we perform Synonym Replacement with a 50% probability and Random Insertion, Random Swap, and Random Deletion with a 20% probability each.
@@ C.1 Hyperparameter Settings
When training the Consistency Motion Generator, we set the patch size to 64 and the mask ratio of the Sparse Causal mask to 0.6 during the training phase, with a batch size of 64 and a learning rate of 4e-5. Similarly, for stability, we implement a 50-step warmup followed by 300 epochs of training.
@@
According to Figure 10, in the inference phase, we utilize model parameters saved at the 30th, 75th, and 70th epochs for the Semantic Decoder, Structural Decoder, and Consistency Motion Generator, respectively. For the generative model, we employ inflated Stable Diffusion V1-5. Given that Stable Diffusion operates on Gaussian-distributed latent space inputs (i.e., $Z_{T}$) for Text-to-Image task, and the distribution of decoded structural information does not align with this, we apply 250 steps of Gaussian smoothing to it, which includes 50 steps of ddim inversion. All experiments are conducted on an A100 80G GPU, with the training phase taking 8 hours and the inference phase taking 12 hours for each dataset.
C.2 Evaluation Metric Implementation
@@
This algorithm performs the N-trial n-way top-1 classification test. We describe our evaluation method in Algorithm 2. For the image-based scenario, we utilize a Vision Transformer (ViT) (Dosovitskiy et al. (2021)) pre-trained on ImageNet as the classifier. For video-based scenario, a pre-trained VideoMAE (Tong et al. (2022)) is employed as the classifier.
@@
Spatiotemporal-level
Considering the input to the video reconstruction task contains substantial noise, there are instances where every pixel of each reconstructed video frame is either zero or noise, which would artificially inflate the CLIP score if used directly. Therefore, we calculate the CLIP score only when the VIFI-CLIP value exceeds the average level; otherwise, we assign a score of 0.
Shuffle test
Since not every test sample is accurately reconstructed, some fail to be reconstructed in terms of semantics, structure, or motion information, and some reconstruction results are entirely noise (failed samples are detailed in Appendix E.6.2). Therefore, it is meaningless to perform the shuffle test on the entire test set; we only conduct the shuffle test on the samples that are successfully reconstructed. For details, please refer to the open-source code.
Appendix D Model Architecture
D.1 Definition of Frequently Used Symbols
The symbols frequently used in this work are defined in Table 7.
Symbol Definition
$X$ Voxel space
$Y$ Pixel space
$Z(k)$ Feature space, $k\in\{semantic,structure,motion\}$
$D(k)$ Feature decoder, $k\in\{semantic,structure,motion\}$
$\Phi(\cdot)$ Encoder of pretrained VQ-VAE
$MLP(\cdot)$ Trainable multilayer perceptron
$Emb(\cdot)$ Linear embedding layer
n Number of voxels
$\mathbf{x}_{i}$ i-th fMRI activity pattern, $\mathbf{x}_{i}\in\mathbb{R}^{1\times n}$
$\mathbf{v}_{i,j}$ Frames of i-th video , $\mathbf{v}_{i,j}\in\mathbb{R}^{1\times 3\times 512\times 512}$, $j\in[1,8]$
$\mathbf{c}_{i}$ i-th text condition of Stable Diffusion, $\mathbf{c}_{i}\in\mathbb{R}^{1\times 20\times 768}$
$\mathbf{f}_{i}$ i-th fMRI embedding, $\mathbf{f}_{i}\in\mathbb{R}^{1\times 512}$
$\mathbf{v}_{i}$ i-th video embedding in CLIP space, $\mathbf{v}_{i}\in\mathbb{R}^{1\times 512}$
$\mathbf{t}_{i}$ i-th text embedding in CLIP space, $\mathbf{t}_{i}\in\mathbb{R}^{1\times 512}$
$\mathbf{W}_{Q},\mathbf{W}_{K},\mathbf{W}_{V}$ Projection matrices of attention mechanisms
$\tau$ Learned temperature parameter
$\alpha,\lambda_{1},\lambda_{2}$ Hyperparameters of semantic loss function
$L$ Number of CMG blocks
$d_{token}$ Dimension of frame tokens
$B$ Batch size
$\|\cdot\|_{2}$ $\mathcal{L}_{2}$-norm operator
$s(\cdot,\cdot)$ Cosine similarity
Table 7: Definition of frequently used symbols
D.2 Consistency Motion Generator
The Temporal Module is tasked with learning the temporal dynamics from the visible frames. Given the severe information redundancy between video frame tokens, we specifically design a Sparse Causal mask. As shown in Figure 4 on the top, during training, the mask is divided into fixed and random components. The fixed mask ensures that each frame cannot access information from subsequent frames, while the random mask maintains sparsity among visible frames, preventing the model from taking shortcuts (Tong et al. (2022)) and accelerating training. During inference, we eliminate the random mask to allow full utilization of information from all preceding frames for predicting future frames.
D.3 Text-to-Image Network Inflation
To leverage pre-trained weights from large-scale image datasets, such as ImageNet, for the pre-training of massive video understanding models, Carreira & Zisserman (2017) pioneered the expansion of filters and pooling kernels in 2D ConvNets into the third dimension to create 3D filters and pooling kernels. This process transforms N×N filters used for images into N×N×N 3D filters, providing a beneficial starting point for 3D video understanding models by utilizing spatial features learned from large-scale image datasets.
@@ Appendix E Additional Experimental Results
E.1 The impact of different spatial-temporal architecture designs.
Due to computational resource constraints, we adopted a design in the CMG architecture that separates temporal and spatial attention modules. To capture motion information between video frames from fMRI, we designed distinct interactions between the temporal and spatial information of fMRI and video data.
For feature interaction, we primarily adopt two strategies:
(1) Cross-attention:
As shown in Eq. (E.1), we treat video representations as Queries and fMRI representations as Keys and Values, using cross-attention to extract useful spatial and temporal information from fMRI.
$\mathbf{z}_{l}=$ $CrossAttention(\mathbf{Q},\mathbf{K},\mathbf{V}),\quad l=1,2,\ldots,L$
$\mathbf{Q}=$ $\mathbf{W_{Q}}^{l}\cdot\mathbf{z}_{l},\quad\mathbf{K}=\mathbf{W_{K}}^{l}\cdot Emb(\mathbf{f}),\quad\mathbf{V}=\mathbf{W_{V}}^{l}\cdot Emb(\mathbf{f}).$ (11)
(2) Adaptive layer normalization:
Inspired by the success of space-time self-attention in video modeling, we modulate the spatial and temporal information of video representations using fMRI representations, as shown in Eq. (12).
$adaLN(\mathbf{z}_{l},\mathbf{f})=Emb(\mathbf{f})_{scale}LayerNorm({z}_{l})+Emb(\mathbf{f})_{shift}.$ (12)
Figure 11: Variants of the Spatial-Temporal Fusion Layer.
Figure 12: Loss curves for variants of the Spatial-Temporal Fusion Layer.
As shown in Figure 11, we designed four different architectures for the Spatial-Temporal Fusion Layer (STFL). Under the same hyperparameters used in Section C, the training/validation loss curves for these architectures are presented in Figure 12.
As shown in Figure 12, regardless of whether cross-attention or adaptive layer normalization is used, any interaction between fMRI representations and temporal information in the video leads to extreme instability in training and difficulty in convergence. We hypothesize that this is due to the significantly lower temporal resolution of fMRI compared to video stimuli, making it difficult to explicitly extract useful temporal information from fMRI. Therefore, in the design of the CMG, we only allow interaction between fMRI representations and spatial information in the video. For temporal information, we use a sparse causal random mask to implicitly learn motion information between frames.
E.2 A Detailed Parameter Sensitivity Analysis on Patch size.
Patch size Semantic-level ↑ Pixel-level ↑ ST-level
2-way-I 2-way-V VIFI-score SSIM PSNR Hue-pcc CLIP-pcc↑ EPE↓
4 0.704 0.792 0.495 0.214 10.257 0.756 0.072 8.473
8 0.698 0.794 0.477 0.148 9.744 0.720 0.043 9.935
16 0.675 0.778 0.481 0.139 9.773 0.761 0.049 7.625
32 0.718 0.798 0.503 0.176 9.330 0.751 0.094 7.671
64 0.812 0.841 0.602 0.321 9.124 0.774 0.425 5.580
Table 8: Parameter sensitivity analysis on Patch size.
We conducted detailed experiments to investigate the sensitivity of video reconstruction to patch size. As shown in Table 8, setting a smaller patch size hinders the model’s learning. This is because a smaller patch size results in a larger number of patches, forcing fine-grained patch-level interactions between fMRI and video representations. However, the low signal-to-noise ratio of fMRI does not support such fine-grained interactions, and the excessive noise may even degrade the quality of the video representations.
E.3 A Detailed Parameter Sensitivity Analysis on $\lambda_{1}$ and $\lambda_{2}$.
Figure 13: Loss curves for variants of $\lambda_{1}$ and $\lambda_{2}$.
Model Semantic-$\uparrow$ Pixel-$\uparrow$ ST-level
2-way-I 2-way-V VIFI-score SSIM PSNR Hue-pcc CLIP-pcc↑ EPE↓
$\lambda_{1}=0.01$, $\lambda_{2}=0.25$ 0.765 0.825 0.581 0.318 10.199 0.780 0.396 6.025
$\lambda_{1}=0.005$, $\lambda_{2}=0.5$ 0.786 0.824 0.591 0.320 9.109 0.776 0.407 5.898
$\lambda_{1}=0.01$, $\lambda_{2}=0.5$ (Ours) 0.812 0.839 0.604 0.319 9.116 0.778 0.413 5.572
Table 9: Parameter sensitivity analysis for variants of $\lambda_{1}$ and $\lambda_{2}$ on subject 1 of CC2017 dataset. 100 sample sets were constructed using the bootstrap method for hypothesis testing. Colors reflect statistical significance (Wilcoxon test for paired samples) compared to the Full Model. $p<0.0001$ (purple); $p<0.01$ (pink); $p<0.05$ (yellow); $p>0.05$ (green).
We conducted a sensitivity analysis on the selection of $\lambda_{1}$ and $\lambda_{2}$ using sub1 from the CC2017 dataset. As shown in Figure 13, when either $\lambda_{1}$ or $\lambda_{2}$ is set to 0, the semantic decider fails to converge, indicating that both the contrastive loss and projection loss play crucial roles in decoding semantic information. In Table 9, we set $\lambda_{1}=0.01$ and $\lambda_{2}=0.5$ to ensure that both loss terms are balanced during optimization. When either $\lambda_{1}$ or $\lambda_{2}$ is adjusted, breaking this balance does not affect the structural level metrics of the reconstruction results, but it does impact the semantic and ST level metrics.
E.4 A Detailed Parameter sensitivity analysis on $\alpha$ and mask ratio.
@@
0.787 0.812 0.584
Table 10: Parameter sensitivity analysis on contrastive learning on subject 1 of CC2017 dataset. Colors reflect statistical significance (Wilcoxon test for paired samples) compared to Our Model. $p<0.0001$ (purple); $p<0.01$ (pink); $p<0.05$ (yellow); $p>0.05$ (green).
@@
Models Pixel-level ↑ ST-level
SSIM PSNR Hue-pcc CLIP-pcc↑ EPE↓
Ratio=0 0.297 9.037 0.758 0.397 5.896
Ratio=0.2 0.276 8.847 0.767 0.382 5.311
Ratio=0.4 0.285 9.045 0.768 0.404 5.375